### v2

# III.1 De Palma-Trevisan-type quantum Wasserstein distance

III.1 De Palma-Trevisan-type quantum Wasserstein distance
It has been found that a quantum Wasserstein distance between quantum states can be defined similarly to Eq. (1),
where the optimization is over separable quantum states [30]. Note that when considering the optimization over separable states, a simpler form has been obtain without transposes [30]. While
its self-distance for is
where is the quantum Fisher information [45, 46, 47, 48, 49]. The definition of was motivated by the fact that the quantum Fisher information is four times the convex roof of the variance [50, 51], and this statement can be reformulated as an optimization over separable states [52], which can be used to estimate the quantum Fisher information based on measurements [53], and to formulate new uncertainty relations [54, 55]. In the more general case of non-unitary dynamics, the quantum Fisher information is not four times the convex roof of the variance any more, but it can be obtained via semidefinite programming [56]. In Eq. (12), the coupling is a separable state [44]
The marginals are
The self-distance for is given as [30]
where is decomposed as in Eq. (16a).
Let us consider the case when is a pure state of any dimension and is an arbitrary density matrix of the same dimension. Then, when computing the various quantum Wasserstein distance measures between and the state in the optimization is constrained to be the tensor product of the two density matrices. Let us define the auxiliary quantity
Hence, for the distance from a pure state [30]
holds. Thus, based on Eq. (18), the distance square is the sum of the self distances for and which equal the sum of variances for those states over two, and another term that is the sum of the squares of the differences in the expectation values of for the two states over two.
The optimization given in Eq. (12) can be carried out numerically exactly using semidefinite programming for the case of qubits, while we can obtain vary good lower bounds for large systems, if we carry out the minimization over quantum states with a positive partial transpose [30]. The Wasserstein distance above is known to fulfill the condition (Eq. (126), Ref. [30])
where for pure states we have
where the conditions for the marginals given in Eq. (16) are fulfilled and is defined in Eq. (18). Note that here we used the notation instead of for shortness, and we will follow this convention in the rest of the paper.
Thus, we can define based on optimizing a quantity that is given for pure states as in Eq. (III.1). This is an advantage, since we can start from a quantity defined for pure states only, and then define a quantity that is also valid for mixed states. However, we also know how to compute it numerically using semidefinite programming based on Eq. (12).
Note that we could also write down an optimization over mixed separable states as
where the separable state is given as
The marginals are
Since in Eq. (12) we optimize an expression linear in expectation values, the optimization in Eq. (III.1) would lead to the same result as in Eq. (III.1).
For for the self-distance
holds, where is decomposed as in Eq. (16a) [30]. Note the important relation between the the Wigner-Yanase skew information and the quantum Fisher information given as
III.2 Swap fidelity and the corresponding quantum Wasserstein distance
Before turning to the swap fidelity, let us obtain a useful relation for Uhlmann-Jozsa fidelity, which we will use later.
The Uhlmann-Jozsa fidelity can be given as [42]
where and are the purifications of and respectively.
We can reformulate Eq. (27) as an optimization over separable states as follows.
Theorem 1. For the Uhlmann-Jozsa fidelity the relation
holds such that the conditions for the marginals given in Eq. (16) are fulfilled.
Proof. Equation (23) in Ref. [57] shows that
where the two states can be decomposed to a mixture of pure states as
and
Clearly, the maximum is taken when are all real and positive [57]. Hence,
also holds. Now, for any we can find such that Eq. (15) and Eq. (16) are fulfilled. In practice, in this case
several can be equal to each other and several can be equal to each other. For clarity, let us see an example. For instance, and Then,
we can choose
Thus, we arrive at Eq. (28).
We can also make the following statement.
Theorem 2. We can give the Bures metric with an optimization over separable states as
Proof. We used that the Bures istance can be given wih the fidelity as [42, 58]
After discussing the Uhlmann-Jozsa fidelity and the Bures metric, let us consider now the swap fidelity based on an optimization over separable states as
where is the set of separable states. We can formulate the following theorem.
Theorem 3. For the maximum over separable states
holds, and the conditions for the marginals given in Eq. (16) are fulfilled. It is instructive to compare Eq. (37) to Eq. (28).
Proof. We have to use that for matrices and the trace of their product can be expressed with their tensor product and the swap operator as
Then, for a separable given in Eq. (15)
holds.
Let us make a comment concerning the relation of this finding to entanglement theory. The relation in Eq. (39) indicates that for any separable state
holds, and any state that violates the inequality in Eq. (40) is entangled. It is also known that any bipartite quantum state in the symmetric subspace violating Eq. (40) has a non-positive partial transpose [59].
Let us now compare the various quantities discussed so far. We find that the inequalities
hold. In particular, the first inequality holds since since for any probabilities and real numbers we have
which is due to the convexity of the function The second inequality holds since on the left-hand side we are optimizing over a subset of physical states, while on the right-hand side we optimize over the set of physical states.
Note that for pure states we have
thus can be given by an optimization over separable states as
such that the conditions for the marginals given in Eq. (16) are fulfilled. The distance based on an optimization over separable states is defined similarly to Eq. (8) as
Then, analogously to the statements about the swap fidelity, we have that
holds such that the conditions for the marginals given in Eq. (16) are fulfilled.
III.3 Modified De Palma-Trevisan-type quantum Wasserstein distance
In Eq. (III.1), we optimized the square of the De Palma-Trevisan-type quantum Wasserstein distance. Next, we examine what happens if we replace in Eq. (III.1) with the square of the modified distance given in Eq. (48).
Let us consider now the quantity defining the Euclidean distance square with the expectation values of as
For simplifying our calculations, we also included the constant in Eq. (47). For pure states, the modified De Palma-Trevisan-type quantum Wasserstein distance is given as
Clearly, if then the above expression is zero. Then, we consider the quantity
such that the conditions for the marginals given in Eq. (16) are fulfilled.
Here the subscript ”decomp” in Eq. (LABEL:eq:GMPC_distance_sep_recursive2) indicates that the quantity is defined with an optimization over decompositions to pure product states.
Based on the definition given in Eq. (LABEL:eq:GMPC_distance_sep_recursive2), we can see that
where is decomposed as in Eq. (16a).
Next, we will find a relevant inequality for the quantity we have just introduced.
Theorem 4. The expression based on a minimization over separable decompositions given in Eq. (LABEL:eq:GMPC_distance_sep_recursive2) is bounded from below with the distance based on the expectation values as
where the optimization is over separable states of the form given in Eq. (23), the marginals fulfill Eq. (24), and is defined in Eq. (47).
Proof. The inequality is true since the minimization is over a larger set on the right-hand side than on the left-hand side. Let us define
Then, let us consider the non-negative quantity
From these, it follows that
The inequality can always be saturated choosing and
Following the idea for the definition of the modified quantum Wasserstein distance in Eq. (11), let us now define an analogous quantity for the
quantum Wasserstein distance as
Clearly, for pure states
holds. Then, we can relate and to each other as follows.
Theorem 5. For the modified quantum Wasserstein distance based on an optimization over separable states
holds. For the case, this means
Proof. Based on Eq. (III.1) and Eq. (21), we can write that
For , the self-distance is given in Eq. (14).
So far we considered an arbitrary set of operators. This makes it possible to choose operators that are important for the problem considered. Moreover, two orthogonal states do not have necessarily a maximal distance. Next, we consider the opposite situation, when we have a full set of traceless pairwise orthogonal operator, i. e., generators, such that (see Ref. [60], and Observation 10 in Ref. [30]). In this case, we have a unitarily invariant distance and two orthogonal pure states do have a maximal distance. In particular, the operators must satisfy (see, e.g., Ref. [60])
For such operators, for pure states for the sum of expectation value squares
holds, while for the sum of the tensor products we have
Then, for the sum of the squares of the differences we have
where is again the swap operator. Hence, we conclude the following.
Theorem 6. If we consider a full set of operators, then the following two quantities are equal to each
We will now show that Eq. (57) can be saturated. We will use that for all pure states [30]
holds, hence the self-distance is independent from the state and it is
Based on these, we find the following.
Theorem 7. The inequality in Eq. (57) is saturated and thus
holds, if we consider a full set of operators fulfilling Eq. (60).
For the proof, see Appendix A. We also present an example with when the inequality is not saturated.
Due to Theorem 7 and the definition of in Eq. (LABEL:eq:GMPC_distance_sep_recursive2), in the special case of having a full set of operators,
holds, while it is not true in general. Here we used that for pure states we have
Next, we relate and to each other. Since the inequality in Eq. (13) holds,
we would expect an analogous relation for the modified quantities. Surprisingly, we find an inequality in the opposite direction.
Theorem 8. The following inequality
holds if is pure.
Proof. The relation given in Eq. (19) holds for the case when is a pure state and is an arbitrary density matrix.
Hence, the difference between the two quantities is just the difference of the two self-distances for over two
where is decomposed as in Eq. (16b), for the first inequality we used Eq. (25), which puts a lower bound on the self-distance with the quantum Fisher information. For the second inequality we used Eq. (26), which is an inequality between the quantum Fisher information over four and the Wigner-Yanase skew information. The self-distances are given in Eqs. (2) and (17). Note that for we have
This finishes our proof.
For using the quantities based on an optimization over separable states defined above to measure how far a quantum state is from another one, it would be important to show that these quantities fulfill the triangle inequality. Let us now examine this in the case of We will take advantage of the fact that it is defined as an optimization of an expression given for two pure states and the optimization is taken over all possible separable decompositions of a bipartite state with given marginals.
Theorem 9. The inequality
where and are density matrices, is fulfilled for the case of a pure
Proof. We know that fulfills an analogous inequality for being pure or both and being pure [29], or one of the three states being pure [28]. From this, it follows trivially that Eq. (72) is true for pure states, since for pure states = holds. However, the triangle inequality can also be proven easily for pure states directly.
From Eq. (LABEL:eq:GMPC_distance_sep_recursive2) follows for an optimal decomposition that
where we used Eq. (42), and and are decomposed as in Eq. (16). Then,
where is decomposed as in Eq. (30), is decomposed as
and is decomposed as
Here, again, the decompositions in Eq. (74) are optimal. Let us now assume that is a pure state. Then,
where is an optimal decomposition obtained from following the ideas in the proof of Theorem 1, and in the second inequality we used that the triangle inequality is valid for pure states.
After considering several possibilities of defining the quantum Wasserstein distance based on an optimization over separable states, let us summarize our findings. In many of the examples in Sec. III, we had the square of the De Palma-Trevisan distance, optimized over separable states
or the square of the modified De Palma-Trevisan distance, optimized over separable states
such that the conditions for the marginals given in Eq. (16) are fulfilled. In the first case, we have a quantity that might have a nonzero self-distance. In the second case, the quantify given by the optimization does not have a self-distance.