### v4

# Several types of quantum Wasserstein distance based on an optimization over separable states

Abstract
We consider several definitions of the quantum Wasserstein distance based on an optimization over general bipartite quantum states with given marginals. Then, we examine the quantities obtained after the optimization is carried out over bipartite separable states instead. We prove that several of these quantities are equal to each other. Thus, we connect several approaches in the literature. We prove the triangle inequality for some of these quantities for the case of one of the three states being pure. As a byproduct, we show that the Uhlmann-Jozsa quantum fidelity can also be written as an optimization over separable states with given marginals. We use this to prove that some of these quantities equal the Uhlmann-Jozsa quantum fidelity for qubits.
I Introduction
Quantum optimal transport has been at the center of attention, as it led to the definition of several new and very useful notions in quantum physics. The first, semi-classical approach of Życzkowski and Słomczyński, has been motivated by applications in quantum chaos [1, 2, 3]. The method of Biane and Voiculescu is related to free probability [4], while the one of Carlen, Maas, Datta, and Rouzé [5, 6, 7, 8, 9] is based on a dynamical interpretation. Caglioti, Golse, Mouhot, and Paul presented an approach based on a static interpretation [10, 11, 12, 13, 14, 15], which has been connected to negative Sobolev norms [16]. Finally De Palma and Trevisan used quantum channels [17], and De Palma, Marvian, Trevisan, and Lloyd defined the quantum earth mover’s distance, i. e., the quantum Wasserstein distance order 1 [18], while Bistroń, Cole, Eckstein, Friedland, and Życzkowski formulated a quantum Wasserstein distance based on an antisymmetric cost function and the swap fidelity [19, 20, 21].
One of the key results of quantum optimal transport is the definition of the quantum Wasserstein distance [1, 2, 3, 10, 11, 12, 13, 17, 18, 14, 15, 22, 23]. It has the often desirable feature that it is not necessarily maximal for two quantum states orthogonal to each other, which is beneficial, for instance, when performing learning on quantum data [24]. Some of the properties of the new quantities are puzzling, yet point to profound relations between seemingly unrelated fields of quantum physics. For instance, the quantum Wasserstein distance order 2 of the quantum state from itself can be nonzero, while in the classical case the self-distance is always zero. In particular, as we have mentioned, the quantum Wasserstein distance has been defined based on a quantum channel formalism for a given set of operators $\{H_{n}\}_{n=1}^{N}$ [17, 25], and it has been shown that the square of the self-distance for a state $\varrho$ is equal to the sum of the Wigner-Yanase skew information $I_{\varrho}(H_{n})$ of the quantum state [26]. Such a quantum Wasserstein distance has been used to examine the distance of various ground state phases in spin systems [27]. It has also been extended to define a $p$-Wasserstein distance [28]. A modified quantum Wasserstein distance has been defined connected to this approach, that does not have a self-distance [17] and a strong evidence suggests that it fulfills the triangle inequalities [29, 28].
Recently, it has been found that if the optimization is taken over separable states then the self-distance equals the quantum Fisher information over four for when a single operator is given, i. e., $N=1$ [30]. This result connects the quantum Wasserstein distance to quantum theory [31, 32, 33] and quantum metrology [34, 35, 36, 37, 38, 39]. In this approach, the formula containing an optimization over bipartite separable states with given marginals provides the distance for mixed states, where the cost function optimized is the distance based on the channel formalism between two pure states [30]. Another definition of the quantum Wasserstein distance based on an optimization over separable states has been considered in which the cost function is the trace distance, and for which the self-distance is zero [40].
Having several possible definitions of the quantum Wasserstein distance, the question arises: Is it possible to connect these approaches to each other, and possibly reduce them to few basic types?
In this paper, we answer this question affirmatively. We consider some of the definitions of the quantum Wasserstein distance, in which the optimization is taken over the set of all quantum states, and examine the quantities that arise when we carry out the optimization over separable states instead. We show that we obtain one of two cases.
(i) In the first case, we obtain an optimization of a cost function given as the distance defined by the quantum channel formalism for pure states [25], which is the case discussed in Ref. [30] and defines a quantity that can have a non-zero self-distance.
(ii) In the second case, we obtain an optimization of a cost function given as the modified distance defined by the quantum channel formalism for pure states [25], which defines a quantity that has a zero self-distance. In the case of a full set of observables $H_{n}$, this quantity equals the definition based on the trace distance [40]. It also equals the distance based on swap fidelity, if we carry out an optimization over separable states rather than over general quantum states [19, 20, 21].
The paper is organized as follows. In Sec. II, we list some of the types of the quantum Wasserstein distance defined with an optimization over general states we consider in this article. In Sec. III, we calculate the above definitions for the case when we carry out the optimization over separable states. We also write down the quantum Fidelity and the Bures distance as an optimization over separable states and relate it to the other distance definitions. In Sec. IV, we consider the case of the quantum $p$-Wasserstein distance.
II Setting the stage
In this section, we list various definitions of the quantum Wasserstein distance. We consider the definitions that involve an optimization over all physical quantum states.
II.1 De Palma-Trevisan-type quantum Wasserstein distance
One of the definitions of the quantum Wasserstein distance has been given by De Palma and Trevisan as [17]
[…]
where $H_{n}$ are Hermitian operators, and $\mathcal{D}$ is the set of physical states. Its self-distance is
$D_{\rm DPT}(\varrho,\varrho)^{2}=\sum_{n}I_{\varrho}(H),$ (2)
where $I$ is the Wigner-Yanase skew information [26] defined as
$I_{\varrho}(H)={\rm Tr}(H^{2}\varrho)-{\rm Tr}(H\sqrt{\varrho}H\sqrt{\varrho}).$ (3)
The advantage of the approach is that there is a transport map corresponding to all $\varrho_{12}$ couplings. Note that this is possible due to the specific form of the definition in Eq. (1), which includes several partial transposes.
II.2 Swap fidelity and the corresponding quantum Wasserstein distance
The swap fidelity is defined as [19, 41, 20]
$F_{S}(\varrho,\sigma)=\max_{\varrho_{12\in\mathcal{D}}}{\rm Tr}(\varrho_{12}S),$ (4)
such that the marginals of $\varrho_{12}$ are $\varrho$ and $\sigma.$ Here $S$ is the swap operator defined as
$S|\Psi\rangle\otimes|\Phi\rangle=|\Phi\rangle\otimes|\Psi\rangle.$ (5)
It is known that [19]
$F(\varrho,\sigma)\leq F_{S}(\varrho,\sigma)\leq\sqrt{F(\varrho,\sigma)},$ (6)
where the Uhlmann-Jozsa fidelity is defined as [42, 43]
$F(\varrho,\sigma)={\rm Tr}\left(\sqrt{\sqrt{\varrho}\sigma\sqrt{\varrho}}\right)^{2}.$ (7)
The corresponding distance is given as [19]
$D_{S}(\varrho,\sigma)^{2}=\frac{1}{2}[1-F_{S}(\varrho,\sigma)],$ (8)
which can be expressed as
$D_{S}(\varrho,\sigma)^{2}=\min_{\varrho_{12}\in\mathcal{D}}{\rm Tr}(\varrho_{12}C_{S})$ (9)
such that the marginals of $\varrho_{12}$ are $\varrho$ and $\sigma,$ and the quantum cost matrix is
[…]
with the property $(C_{S})^{2}=C_{S}.$
II.3 Modified De Palma-Trevisan-type quantum Wasserstein distance
A modified De Palma-Trevisan distance has been considered that does not have a nonzero self-distance [25]
$\begin{split}\tilde{D}_{\rm DPT}(\varrho,\sigma)^{2}&=D_{\rm DPT}(\varrho,\sigma)^{2}\\ &-[D_{\rm DPT}(\varrho,\varrho)^{2}+D_{\rm DPT}(\sigma,\sigma)^{2}]/2.\end{split}$ (11)
There is strong evidence that the expression fulfills the triangle inequality [29, 28].
III Wasserstein distance based on an optimization over separable quantum states
In this section, we consider the definitions given in Sec. II, and ask what happens if we optimize over separable quantum states instead of general quantum states. Separble quantum states are just mixtures of product states [44]. Any quantum state that is not separable is called entangled. (For a review on entanglement theory, see, e. g., Refs. [31, 32, 33]).
III.1 De Palma-Trevisan-type quantum Wasserstein distance
It has been found that a quantum Wasserstein distance between quantum states can be defined similarly to Eq. (1),
[…]
where the optimization is over separable quantum states [30]. Note that when considering the optimization over separable states, a simpler form has been obtained without transposes [30]. While
$D_{\rm DPT,sep}(\varrho,\sigma)^{2}\geq D_{\rm DPT}(\varrho,\sigma)^{2},$ (13)
its self-distance for $N=1$ is
$D_{\rm DPT,sep}(\varrho,\varrho)^{2}=\frac{1}{4}F_{Q}[\varrho,H_{1}],$ (14)
where $F_{Q}$ is the quantum Fisher information [45, 46, 47, 48, 49]. The definition of $D_{\rm DPT,sep}(\varrho,\sigma)^{2}$ was motivated by the fact that the quantum Fisher information is four times the convex roof of the variance [50, 51], and this statement can be reformulated as an optimization over separable states [52], which can be used to estimate the quantum Fisher information based on measurements [53], and to formulate new uncertainty relations [54, 55]. In the more general case of non-unitary dynamics, the quantum Fisher information is not four times the convex roof of the variance any more, but it can be obtained via semidefinite programming [56]. In Eq. (12), the coupling $\varrho_{12}$ is a separable state [44]
$\varrho_{12}=\sum_{k}p_{k}|\Psi_{k}\rangle\langle\Psi_{k}|\otimes|\Phi_{k}\rangle\langle\Phi_{k}|.$ (15)
The marginals are
$\varrho$ $=\sum_{k}p_{k}|\Psi_{k}\rangle\langle\Psi_{k}|,$ (16a)
$\sigma$ $=\sum_{k}p_{k}|\Phi_{k}\rangle\langle\Phi_{k}|.$ (16b)
The self-distance for $N>1$ is given as [30]
$D_{\rm DPT,sep}(\varrho,\varrho)^{2}=\min_{\{p_{k},|\Psi_{k}\rangle\}}\sum_{n=1}^{N}\sum_{k}p_{k}(\Delta H_{n})^{2}_{\Psi_{k}},$ (17)
where $\varrho$ is decomposed as in Eq. (16a).
Let us consider the case when $\varrho=|\Psi\rangle\langle\Psi|$ is a pure state of any dimension and $\sigma$ is an arbitrary density matrix of the same dimension. Then, when computing the various quantum Wasserstein distance measures between $\varrho$ and $\sigma,$ the state $\varrho_{12}$ in the optimization is constrained to be the tensor product of the two density matrices. Let us define the auxiliary quantity
$\begin{split}&\Delta(\varrho,\sigma)^{2}=\frac{1}{2}\sum_{n=1}^{N}\bigg{[}(\Delta H_{n})^{2}_{\varrho}+(\Delta H_{n})^{2}_{\sigma}\\ &\quad\quad\quad\quad\quad\quad+(\langle{H_{n}}\rangle_{\varrho}-\langle{H_{n}}\rangle_{\sigma})^{2}\bigg{]}.\end{split}$ (18)
Hence, for the distance from a pure state [30]
$\begin{split}D_{\rm DPT}(|\Psi\rangle\langle\Psi|,\sigma)^{2}&=D_{\rm DPT,sep}(|\Psi\rangle\langle\Psi|,\sigma)^{2}\\ &=\Delta(|\Psi\rangle\langle\Psi|,\sigma)^{2}\end{split}$ (19)
holds. Thus, based on Eq. (18), the distance squared is the sum of the self-distances for $\varrho$ and $\sigma,$ which equal the sum of variances for those states over two, and another term that is the sum of the squares of the differences in the expectation values of $H_{n}$ for the two states over two.
The optimization given in Eq. (12) can be carried out numerically exactly using semidefinite programming for the case of qubits, while we can obtain vary good lower bounds for large systems, if we carry out the minimization over quantum states with a positive partial transpose [30]. The Wasserstein distance above is known to fulfill the condition (see Eq. (126) in Ref. [30])
$D_{\rm DPT,sep}(\varrho,\sigma)^{2}\quad\quad\quad$
$=\min_{\{p_{k},|\Psi_{k}\rangle,|\Phi_{k}\rangle\}}\;$ $\sum_{k}p_{k}D_{\rm DPT}(|\Psi_{k}\rangle,|\Phi_{k}\rangle)^{2},$ (20)
where for pure states we have
$\begin{split}D_{\rm DPT}(|\Psi\rangle,|\Phi\rangle)^{2}&=D_{\rm DPT,sep}(|\Psi\rangle,|\Phi\rangle)^{2}\\ &=\Delta(|\Psi\rangle,|\Phi\rangle)^{2},\end{split}$ (21)
where the conditions for the marginals given in Eq. (16) are fulfilled and $\Delta(\varrho,\sigma)^{2}$ is defined in Eq. (18). Note that here we used the notation $D_{\rm DPT}(|\Psi\rangle,|\Phi\rangle)^{2}$ instead of $D_{\rm DPT}(|\Psi\rangle\langle\Psi|,|\Phi\rangle\langle\Phi|)^{2}$ for shortness, and we will follow this convention in the rest of the paper.
Thus, we can define $D_{\rm DPT,sep}(\varrho,\sigma)^{2}$ based on optimizing a quantity that is given for pure states as in Eq. (III.1). This is an advantage, since we can start from a quantity defined for pure states only, and then define a quantity that is also valid for mixed states. However, we also know how to compute it numerically using semidefinite programming based on Eq. (12).
An upper bound can be found based on an optimization only over one of the two states as
$D_{\rm DPT,sep}(\varrho,\sigma)^{2}\leq\min_{\{p_{k},\Psi_{k}\}}\sum_{k}p_{k}D_{\rm DPT,sep}(|\Psi_{k}\rangle,\sigma)^{2}.$ (22)
where for $\varrho$ Eq. (16a) holds.
Note that we could also write down an optimization over mixed separable states as
$D_{\rm DPT,sep}(\varrho,\sigma)^{2}\quad\quad\quad$
$=\min_{\{p_{k},\varrho_{k},\sigma_{k}\}}\;$ $\sum_{k}p_{k}D_{\rm DPT}(\varrho_{k},\sigma_{k})^{2},$ (23)
where the separable state is given as
$\varrho_{12}=\sum_{k}p_{k}\varrho_{k}\otimes\sigma_{k}.$ (24)
The marginals are
$\begin{split}\varrho&=\sum_{k}p_{k}\varrho_{k},\\ \sigma&=\sum_{k}p_{k}\sigma_{k}.\end{split}$ (25)
Since in Eq. (12) we optimize an expression linear in expectation values, the optimization in Eq. (III.1) would lead to the same result as in Eq. (III.1).
For $N>1,$ for the self-distance
$\frac{1}{4}\sum_{n=1}^{N}F_{Q}[\varrho,H_{n}]$
$\quad\quad\leq D_{\rm DPT,sep}(\varrho,\varrho)^{2}=\min_{\{p_{k},|\Psi_{k}\rangle\}}$ $\sum_{n=1}^{N}\sum_{k}p_{k}(\Delta H_{n})^{2}_{\Psi_{k}}$
$\quad\quad\leq\sum_{n=1}^{N}(\Delta H_{n})^{2}$ (26)
holds, where $\varrho$ is decomposed as in Eq. (16a) [30]. Note the important relation between the the Wigner-Yanase skew information and the quantum Fisher information given as
$I_{\varrho}(H)\leq\frac{1}{4}F_{Q}[\varrho,H].$ (27)
The quantum Wasserstein distance can be bounded from below with the self-distances as
$D_{\rm DPT,sep}(\varrho,\sigma)^{2}\geq\frac{1}{2}[D_{\rm DPT,sep}(\varrho,\varrho)^{2}+D_{\rm DPT,sep}(\sigma,\sigma)^{2}],$ (28)
see Eq. (133) in Ref. [30].
Finally, note that even for $D_{\rm DPT,sep}(\varrho,\sigma)^{2}$ there is a transport map corresponding to all separable $\varrho_{12}$ couplings, see Eq. (83) in Ref. [30].
III.2 Swap fidelity and the corresponding quantum Wasserstein distance
Before turning to the swap fidelity, let us obtain a useful relation for the Uhlmann-Jozsa fidelity, which we will use later.
The Uhlmann-Jozsa fidelity can be given as [42]
$F(\varrho,\sigma)=\max_{\Psi_{\varrho},\Phi_{\sigma}}|\langle\Psi_{\varrho}|\Phi_{\sigma}\rangle|^{2},$ (29)
where $|\Psi_{\varrho}\rangle$ and $|\Phi_{\sigma}\rangle$ are the purifications of $\varrho$ and $\sigma,$ respectively.
We can reformulate Eq. (29) as an optimization over separable states as follows.
Theorem 1. For the Uhlmann-Jozsa fidelity the relation
$F(\varrho,\sigma)=\left(\max_{\{p_{k},\Psi_{k},\Phi_{k}\}}\sum_{k}p_{k}|\langle\Psi_{k}|\Phi_{k}\rangle|\right)^{2},$ (30)
holds such that the conditions for the marginals given in Eq. (16) are fulfilled.
Proof. Equation (23) in Ref. [57] shows that
$F(\varrho,\sigma)=\left|\max_{\{\tilde{p}_{k},\tilde{q}_{k},\tilde{\Psi}_{k},\tilde{\Phi}_{k}\}}\sum_{k}\sqrt{\tilde{p}_{k}\tilde{q}_{k}}\langle\tilde{\Psi}_{k}|\tilde{\Phi}_{k}\rangle\right|^{2},$ (31)
where the two states can be decomposed to a mixture of pure states as
$\varrho=\sum_{k}\tilde{p}_{k}|\tilde{\Psi}_{k}\rangle\langle\tilde{\Psi}_{k}|$ (32)
and
$\sigma=\sum_{k}\tilde{q}_{k}|\tilde{\Phi}_{k}\rangle\langle\tilde{\Phi}_{k}|.$ (33)
Clearly, the maximum is taken when $\langle\tilde{\Psi}_{k}|\tilde{\Phi}_{k}\rangle$ are all real and positive [57]. Hence,
$F(\varrho,\sigma)=\left(\max_{\{\tilde{p}_{k},\tilde{q}_{k},\tilde{\Psi}_{k},\tilde{\Phi}_{k}\}}\sum_{k}\sqrt{\tilde{p}_{k}\tilde{q}_{k}}\left|\langle\tilde{\Psi}_{k}|\tilde{\Phi}_{k}\rangle\right|\right)^{2}$ (34)
also holds. Now, for any $\{\tilde{p}_{k},\tilde{q}_{k},\tilde{\Phi}_{k},\tilde{\Psi}_{k}\}$ we can find $\{p_{k},\Psi_{k},\Phi_{k}\}$ such that Eq. (15) and Eq. (16) are fulfilled. In practice, in this case several $|\Psi_{k}\rangle$ can be equal to each other and several $|\Phi_{k}\rangle$ can be equal to each other. For clarity, let us see an example. For instance, $\{p_{k}\}=\{0.5,0.5\}$ and $\{q_{k}\}=\{0.75,0.25\}.$ Then, we can choose
$\begin{split}&p_{1}=0.50,|\Psi_{1}\rangle=|\tilde{\Psi}_{1}\rangle,|\Phi_{1}\rangle=|\tilde{\Phi}_{1}\rangle,\\ &p_{2}=0.25,|\Psi_{2}\rangle=|\tilde{\Psi}_{2}\rangle,|\Phi_{2}\rangle=|\tilde{\Phi}_{1}\rangle,\\ &p_{3}=0.25,|\Psi_{3}\rangle=|\tilde{\Psi}_{2}\rangle,|\Phi_{3}\rangle=|\tilde{\Phi}_{2}\rangle.\end{split}$ (35)
[…]
We can also make the following statement.
Theorem 2. We can give the Bures metric with an optimization over separable states as
$D_{\rm Bures}(\varrho,\sigma)^{2}=2\min_{\{p_{k},\Psi_{k},\Phi_{k}\}}\sum_{k}p_{k}(1-\left|\langle\Psi_{k}|\Phi_{k}\rangle\right|).$ (36)
Proof. We used that the Bures distance can be given with the fidelity as [42, 58]
$D_{\rm Bures}(\varrho,\sigma)^{2}=2\left[1-\sqrt{F(\varrho,\sigma)}\right].$ (37)
[…]
After discussing the Uhlmann-Jozsa fidelity and the Bures metric, let us consider now the swap fidelity based on an optimization over separable states as
$F_{S,{\rm sep}}(\varrho,\sigma)=\max_{\varrho_{12}\in\mathrm{Sep}}{\rm Tr}(\varrho_{12}S),$ (38)
where $\mathrm{Sep}$ is the set of separable states. We can formulate the following theorem.
Theorem 3. For the maximum over separable states
$F_{S,{\rm sep}}(\varrho,\sigma)=\max_{\{p_{k},\Phi_{k},\Psi_{k}\}}\sum_{k}p_{k}|\langle\Phi_{k}|\Psi_{k}\rangle|^{2}$ (39)
holds, and the conditions for the marginals given in Eq. (16) are fulfilled. It is instructive to compare Eq. (39) to Eq. (30).
Proof. We have to use that for matrices $A$ and $B$ the trace of their product can be expressed with their tensor product and the swap operator as
${\rm Tr}(AB)={\rm Tr}[(A\otimes B)S].$ (40)
Then, for a separable $\varrho_{12}$ given in Eq. (15)
${\rm Tr}(\varrho_{12}S)=\sum_{k}p_{k}|\langle\Phi_{k}|\Psi_{k}\rangle|^{2}$ (41)
[…]
Let us make a comment concerning the relation of this finding to entanglement theory. The relation in Eq. (41) indicates that for any separable state $\varrho_{12}$
${\rm Tr}(\varrho_{12}S)\geq 0,$ (42)
holds, and any state that violates the inequality in Eq. (42) is entangled. It is also known that any bipartite quantum state in the symmetric subspace violating Eq. (42) has a non-positive partial transpose [59].
Let us now compare the various quantities discussed so far. We find that the inequalities
$F(\varrho,\sigma)\leq F_{S,{\rm sep}}(\varrho,\sigma)\leq F_{S}(\varrho,\sigma)$ (43)
hold. In particular, the first inequality holds since since for any $p_{k}$ probabilities and $x_{k}$ real numbers we have
$\sum_{k}p_{k}x_{k}^{2}\geq\left(\sum_{k}p_{k}x_{k}\right)^{2},$ (44)
which is due to the convexity of the function $f(x)=x^{2}.$ The second inequality holds since on the left-hand side we are maximizing a quantity over a subset of physical states, while on the right-hand side are maximizing over the entire set of physical states.
Note that for pure states we have
$F_{S}(|\Psi\rangle,|\Phi\rangle)=|\langle\Phi|\Psi\rangle|^{2},$ (45)
thus $F_{S,{\rm sep}}(\varrho,\sigma)$ can be given by an optimization over separable states as
$F_{S,{\rm sep}}(\varrho,\sigma)=\max_{\{p_{k},\Phi_{k},\Psi_{k}\}}\sum_{k}p_{k}F_{S}(|\Psi_{k}\rangle,|\Phi_{k}\rangle)$ (46)
such that the conditions for the marginals given in Eq. (16) are fulfilled. The distance based on an optimization over separable states is defined similarly to Eq. (8) as
$D_{S,{\rm sep}}(\varrho,\sigma)^{2}=\frac{1}{2}[1-F_{S,{\rm sep}}(\varrho,\sigma)].$ (47)
Then, analogously to the statements about the swap fidelity, we have that
$D_{S,{\rm sep}}(\varrho,\sigma)^{2}=\min_{\{p_{k},\Phi_{k},\Psi_{k}\}}\sum_{k}p_{k}D_{S}(|\Psi_{k}\rangle,|\Phi_{k}\rangle)^{2}$ (48)
holds such that the conditions for the marginals given in Eq. (16) are fulfilled.
Moreover, for a pure state $\varrho=|\Psi\rangle\langle\Psi|$ and a mixed state $\sigma$ we have all the inequalities in Eq. (43) saturated and hence
$F(|\Psi\rangle\langle\Psi|,\sigma)=F_{S,{\rm sep}}(|\Psi\rangle\langle\Psi|,\sigma)=F_{S}(|\Psi\rangle\langle\Psi|,\sigma).$ (49)
Next, we see another special case.
Theorem 4. For qubits, the swap fidelity computed based on an optimization over separable states and the Uhlmann-Jozsa fidelity are equal to each other
$F_{S,{\rm sep}}(\varrho,\sigma)=F(\varrho,\sigma).$ (50)
Proof. The inequality in Eq. (44) is saturated when all $x_{k}$ are equal to each other. Hence, the first inequality in Eq. (43) is saturated if there is an optimal $\{p_{k},\Psi_{k},\Phi_{k}\}$ decomposition such that
$|\langle\Psi_{k}|\Phi_{k}\rangle|^{2}=F_{S,{\rm sep}}(\varrho,\sigma)=F(\varrho,\sigma)$ (51)
for all $k.$
There is always such a decomposition for qubits. Consider the polar decomposition $\sqrt{\varrho}\sqrt{\sigma}=|\sqrt{\varrho}\sqrt{\sigma}|U,$ where $U$ is unitary and take the eigenvalue equation of $|\sqrt{\varrho}\sqrt{\sigma}|.$
$\begin{split}|\sqrt{\varrho}\sqrt{\sigma}||u_{1}\rangle&=\alpha_{1}|u_{1}\rangle,\\ |\sqrt{\varrho}\sqrt{\sigma}||u_{2}\rangle&=\alpha_{2}|u_{2}\rangle,\end{split}$ (52)
where $\alpha_{1},\alpha_{2}\geq 0$ and $|u_{1}\rangle,|u_{2}\rangle$ form an orthonormal basis. Let us define the vectors as
$\begin{split}|v_{1}\rangle&=\frac{1}{\sqrt{2}}(|u_{1}\rangle+i|u_{2}\rangle),\\ |v_{2}\rangle&=\frac{1}{\sqrt{2}}(|u_{1}\rangle-i|u_{2}\rangle),\end{split}$ (53)
they also form an orthonormal basis. Then for the vectors
$\begin{split}|\Psi_{k}\rangle&=\sqrt{2}\sqrt{\varrho}|v_{k}\rangle,\\ |\Phi_{k}\rangle&=\sqrt{2}\sqrt{\sigma}U^{*}|v_{k}\rangle\end{split}$ (54)
for $k=1,2,$ we have
$\begin{split}\varrho&=\frac{1}{2}|\Psi_{1}\rangle\langle\Psi_{1}|+\frac{1}{2}|\Psi_{2}\rangle\langle\Psi_{2}|,\\ \sigma&=\frac{1}{2}|\Phi_{1}\rangle\langle\Phi_{1}|+\frac{1}{2}|\Phi_{2}\rangle\langle\Phi_{2}|.\\ \end{split}$ (55)
Then it is easy to check that
$\begin{split}|\langle\Psi_{k}|\Phi_{k}\rangle|&=2|\langle v_{k}|\sqrt{\varrho}\sqrt{\sigma}U^{*}|v_{k}\rangle|=2|\langle v_{k}||\sqrt{\varrho}\sqrt{\sigma}||v_{k}\rangle|\\ &=\alpha_{1}+\alpha_{2}={\rm Tr}(|\sqrt{\varrho}\sqrt{\sigma}|)=\sqrt{F(\varrho,\sigma)}.\end{split}$ (56)
for $k=1,2.$$\blacksquare$
III.3 Modified De Palma-Trevisan-type quantum Wasserstein distance
In Eq. (III.1), we optimized the square of the De Palma-Trevisan-type quantum Wasserstein distance. Next, we examine what happens if we replace $D_{\rm DPT}(|\Psi_{k}\rangle,|\Phi_{k}\rangle)^{2}$ in Eq. (III.1) with the square of the modified distance $\tilde{D}_{\rm DPT}(|\Psi_{k}\rangle,|\Phi_{k}\rangle)^{2}$ given in Eq. (58). Let us consider now the quantity defining the Euclidean distance squared with the expectation values of $H_{n}$ as
$E(\varrho,\sigma)^{2}=\frac{1}{2}\sum_{n=1}^{N}(\langle{H_{n}}\rangle_{\varrho}-\langle{H_{n}}\rangle_{\sigma})^{2}.$ (57)
For simplifying our calculations, we also included the constant $1/2$ in Eq. (57). For pure states, the modified De Palma-Trevisan-type quantum Wasserstein distance is given as
$\tilde{D}_{\rm DPT}(|\Psi\rangle,|\Phi\rangle)^{2}=E(|\Psi\rangle,|\Phi\rangle)^{2}.$ (58)
Clearly, if $|\Psi\rangle=|\Phi\rangle$ then the above expression is zero. Then, we consider the quantity
$\begin{split}&\tilde{D}_{\rm DPT,decomp}(\varrho,\sigma)^{2}\\ &\quad=\min_{\{p_{k},|\Psi_{k}\rangle,|\Phi_{k}\rangle\}}\sum_{k}p_{k}\tilde{D}_{\rm DPT}(|\Psi_{k}\rangle,|\Phi_{k}\rangle)^{2},\end{split}$ (59)
such that the conditions for the marginals given in Eq. (16) are fulfilled. Here the subscript ”decomp” in Eq. (LABEL:eq:GMPC_distance_sep_recursive2) indicates that the quantity is defined with an optimization over decompositions to pure product states. Based on the definition given in Eq. (LABEL:eq:GMPC_distance_sep_recursive2), we can see that
$\begin{split}&\tilde{D}_{\rm DPT,decomp}(\varrho,\sigma)^{2}\\ &\quad\leq\min_{\{p_{k},|\Psi_{k}\rangle\}}\sum_{k}p_{k}\tilde{D}_{\rm DPT,decomp}(|\Psi_{k}\rangle,\sigma)^{2},\end{split}$ (60)
where $\varrho$ is decomposed as in Eq. (16a).
Next, we will find a relevant inequality for the quantity we have just introduced.
Theorem 5. The expression based on a minimization over separable decompositions given in Eq. (LABEL:eq:GMPC_distance_sep_recursive2) is bounded from below with the distance based on the expectation values as
$\tilde{D}_{\rm DPT,decomp}(\varrho,\sigma)^{2}$ $\geq\min_{\{p_{k},\varrho_{k},\sigma_{k}\}}\;\sum_{k}p_{k}E(\varrho_{k},\sigma_{k})^{2}$
$=E(\varrho,\sigma)^{2},$ (61)
where the optimization is over separable states of the form given in Eq. (24), the marginals fulfill Eq. (25), and $E(\varrho,\sigma)^{2}$ is defined in Eq. (57).
Proof. The inequality is true since the minimization is over a larger set on the right-hand side than on the left-hand side. Let us define
$\begin{split}Q_{n,k}&=\langle H_{n}\rangle_{\varrho_{k}}-\langle H_{n}\rangle_{\sigma_{k}},\\ W_{n}&=\langle H_{n}\rangle_{\varrho}-\langle H_{n}\rangle_{\sigma}.\end{split}$ (62)
Then, let us consider the non-negative quantity
$Z^{2}$ $=\sum_{n=1}^{N}\sum_{k}p_{k}(Q_{n,k}-W_{n})^{2}$
$=\sum_{n=1}^{N}\left(\sum_{k}p_{k}Q_{n,k}^{2}\right)+W_{n}^{2}-2\left(\sum_{k}p_{k}Q_{n,k}\right)W_{n}$
$=\sum_{n=1}^{N}\sum_{k}p_{k}Q_{n,k}^{2}-W_{n}^{2}.$ (63)
From these, it follows that
$\sum_{n=1}^{N}\sum_{k}p_{k}Q_{n,k}^{2}=Z^{2}+W_{n}^{2}\geq W_{n}^{2}.$ (64)
The inequality can always be saturated choosing $p_{1}=1,\varrho_{1}=\varrho,$ and $\sigma_{1}=\sigma.$$\blacksquare$
Following the idea for the definition of the modified quantum Wasserstein distance in Eq. (11), let us now define an analogous quantity for the quantum Wasserstein distance $D_{\rm DPT,sep}(\varrho,\sigma)$ as
$\begin{split}&\tilde{D}_{\rm DPT,sep}(\varrho,\sigma)^{2}=D_{\rm DPT,sep}(\varrho,\sigma)^{2}\\ &\quad-[D_{\rm DPT,sep}(\varrho,\varrho)^{2}+D_{\rm DPT,sep}(\sigma,\sigma)^{2}]/2.\end{split}$ (65)
This is a non-negative quantity due to the inequality in Eq. (28). Clearly, for pure states
$\tilde{D}_{\rm DPT,decomp}(\varrho,\sigma)^{2}=\tilde{D}_{\rm DPT}(\varrho,\sigma)^{2}$ (66)
holds. Then, we can relate $\tilde{D}_{\rm DPT,decomp}(\varrho,\sigma)^{2}$ and $\tilde{D}_{\rm DPT,sep}(\varrho,\sigma)^{2}$ to each other as follows.
Theorem 6. For the modified quantum Wasserstein distance based on an optimization over separable states
$\tilde{D}_{\rm DPT,decomp}(\varrho,\sigma)^{2}\leq\tilde{D}_{\rm DPT,sep}(\varrho,\sigma)^{2}$ (67)
holds. For the $N=1$ case, this means
$\begin{split}&\tilde{D}_{\rm DPT,decomp}(\varrho,\sigma)^{2}\leq D_{\rm DPT,sep}(\varrho,\sigma)^{2}\\ &\quad-(F_{Q}[\varrho,H_{1}]+F_{Q}[\sigma,H_{1}])/8.\end{split}$ (68)
Proof. Based on Eq. (III.1) and Eq. (21), we can write that
$\begin{split}&D_{\rm DPT,sep}(\varrho,\sigma)^{2}\quad\quad\quad\\ &\quad=\min_{\{p_{k},|\Psi_{k}\rangle,|\Phi_{k}\rangle\}}\;\sum_{k}p_{k}D_{\rm DPT}(|\Psi_{k}\rangle,|\Phi_{k}\rangle)^{2}\\ &\quad\geq[D_{\rm DPT,sep}(\varrho,\varrho)^{2}+D_{\rm DPT,sep}(\sigma,\sigma)^{2}]/2\\ &\quad\quad+\min_{\{p_{k},|\Psi_{k}\rangle,|\Phi_{k}\rangle\}}\;\sum_{k}p_{k}\tilde{D}_{\rm DPT}(|\Psi_{k}\rangle,|\Phi_{k}\rangle)^{2}.\end{split}$ (69)
[…]
So far we considered an arbitrary set of $H_{n}$ operators. This makes it possible to choose operators that are important for the problem considered. Moreover, two orthogonal states do not have necessarily a maximal distance. Next, we consider the opposite situation, when we have a full set of $(d^{2}-1)$ traceless pairwise orthogonal $H_{n}$ operator, i. e., $SU(d)$ generators, such that (see Ref. [60], and Observation 10 in Ref. [30]). In this case, we have a unitarily invariant distance and two orthogonal pure states do have a maximal distance. In particular, the operators must satisfy (see, e.g., Ref. [60])
${\rm Tr}(H_{n}H_{n^{\prime}})=2\delta_{nn^{\prime}}.$ (70)
For such operators, for pure states for the sum of expectation value squares
$\sum_{n=1}^{d^{2}-1}\langle{H_{n}}\rangle^{2}=2(1-1/d)$ (71)
holds, while for the sum of the tensor products we have
[…]
Then, for the sum of the squares of the differences we have
$\frac{1}{2}\sum_{n=1}^{d^{2}-1}(\langle{H_{n}}\rangle_{\Psi}-\langle{H_{n}}\rangle_{\Phi})^{2}$
$\quad\quad=\frac{1}{2}\sum_{n=1}^{d^{2}-1}\langle{H_{n}}\rangle_{\Psi}^{2}+\langle{H_{n}}\rangle_{\Phi}^{2}-\sum_{n=1}^{d^{2}-1}\langle{H_{n}\otimes H_{n}}\rangle_{\Psi\otimes\Phi}$
$\quad\quad=2(1-\langle{S}\rangle_{\Psi\otimes\Phi}),$ (73)
where $S$ is again the swap operator. Hence, we conclude the following.
Theorem 7. If we consider a full set of $H_{n}$ operators, then the following two quantities are equal to each
$\begin{split}\tilde{D}_{\rm DPT,decomp}(\varrho,\sigma)^{2}&=2\left[1-F_{S,\rm sep}(\varrho,\sigma)^{2}\right]\\ &=4D_{S,\rm sep}(\varrho,\sigma)^{2}.\end{split}$ (74)
We will now show that Eq. (67) can be saturated. We will use that for all pure states [30]
$\sum_{n=1}^{d^{2}-1}(\Delta H_{n})^{2}=2(d-1)$ (75)
holds, hence the self-distance is independent from the state and it is
$D_{\rm DPT,sep}(\varrho,\varrho)^{2}=2(d-1).$ (76)
Based on these, we find the following.
Theorem 8. The inequality in Eq. (67) is saturated and thus
$\tilde{D}_{\rm DPT,decomp}(\varrho,\sigma)^{2}=\tilde{D}_{\rm DPT,sep}(\varrho,\sigma)^{2}$ (77)
holds, if we consider a full set of $H_{n}$ operators fulfilling Eq. (70).
For the proof, see Appendix A. We also present an example with $N=1,$ when the inequality is not saturated.
Due to Theorem 8 and the definition of $\tilde{D}_{\rm DPT,decomp}(\varrho,\sigma)^{2}$ in Eq. (LABEL:eq:GMPC_distance_sep_recursive2), in the special case of having a full set of $H_{n}$ operators,
$\begin{split}&\tilde{D}_{\rm DPT,sep}(\varrho,\sigma)^{2}\\ &\quad=\min_{\{p_{k},|\Psi_{k}\rangle,|\Phi_{k}\rangle\}}\sum_{k}p_{k}\tilde{D}_{\rm DPT,sep}(|\Psi_{k}\rangle,|\Phi_{k}\rangle)^{2}\end{split}$ (78)
holds, while it is not true in general. Here we used that for pure states we have $D_{\rm DPT,sep}(\varrho,\sigma)^{2}=D_{\rm DPT}(\varrho,\sigma)^{2}.$
Due to Theorem 4, Theorem 7 and Theorem 8, for the qubit case, if we choose $H_{1}=\sigma_{x},H_{2}=\sigma_{y},H_{3}=\sigma_{z},$ we have
$\tilde{D}_{\rm DPT,sep}(\varrho,\sigma)^{2}=D_{\rm DPT,sep}(\varrho,\sigma)^{2}-2=2-2F(\varrho,\sigma).$ (79)
Thus, in this case we know $\tilde{D}_{\rm DPT,sep}(\varrho,\sigma)^{2}$ with an explict formula. Here we used the self-distance given in Eq. (76) to express $\tilde{D}_{\rm DPT,sep}(\varrho,\sigma)^{2}$ with $D_{\rm DPT,sep}(\varrho,\sigma)^{2}.$
Let us again consider the case of having a full set of $(d^{2}-1)$ operators $H_{n}$ fulfilling Eq. (70). Let us assume that $\varrho=|\Psi\rangle\langle\Psi|$ is pure. Then, we have
$\begin{split}&\tilde{D}_{\rm DPT,decomp}(|\Psi\rangle\langle\Psi|,\sigma)^{2}=\tilde{D}_{\rm DPT,sep}(|\Psi\rangle\langle\Psi|,\sigma)^{2}\\ &\quad\quad=2[1-F_{\rm S,sep}(|\Psi\rangle\langle\Psi|,\sigma)]=2[1-F(|\Psi\rangle\langle\Psi|,\sigma)]\\ &\quad\quad=2[1-{\rm Tr}(|\Psi\rangle\langle\Psi|\sigma)],\end{split}$ (80)
where we used Eqs. (49), (74), and (77).
Next, we relate $\tilde{D}_{\rm DPT,sep}(\varrho,\sigma)^{2}$ and $\tilde{D}_{\rm DPT}(\varrho,\sigma)^{2}$ to each other. Since the inequality in Eq. (13) holds, we would expect an analogous relation for the modified quantities. Unexpectedly, we find an inequality in the opposite direction.
Theorem 9. The following inequality
$\tilde{D}_{\rm DPT,sep}(\varrho,\sigma)^{2}\leq\tilde{D}_{\rm DPT}(\varrho,\sigma)^{2}$ (81)
holds if $\varrho$ is pure.
Proof. The relation given in Eq. (19) holds for the case when $\varrho=|\Psi\rangle\langle\Psi|$ is a pure state and $\sigma$ is an arbitrary density matrix. Hence, the difference between the two quantities is just the difference of the two self-distances for $\sigma$ over two
$\begin{split}&\tilde{D}_{\rm DPT}(\varrho,\sigma)^{2}-\tilde{D}_{\rm DPT,sep}(\varrho,\sigma)^{2}\\ &=\frac{1}{2}\left[\min_{\{p_{k},|\Phi_{k}\rangle\}}\sum_{n=1}^{N}\sum_{k}p_{k}(\Delta H_{n})^{2}_{\Phi_{k}}-\sum_{n=1}^{N}I_{\sigma}(H_{n})\right]\\ &\geq\frac{1}{2}\left[\sum_{n=1}^{N}\frac{1}{4}F_{Q}[\sigma,H_{n}]-\sum_{n=1}^{N}I_{\sigma}(H_{n})\right]\\ &\geq 0,\end{split}$ (82)
where $\sigma$ is decomposed as in Eq. (16b), for the first inequality we used Eq. (26), which puts a lower bound on the self-distance with the quantum Fisher information. For the second inequality we used Eq. (27), which is an inequality between the quantum Fisher information over four and the Wigner-Yanase skew information. The self-distances are given in Eqs. (2) and (17). Note that for $N=1,$ we have
$\begin{split}&\tilde{D}_{\rm DPT}(\varrho,\sigma)^{2}-\tilde{D}_{\rm DPT,sep}(\varrho,\sigma)^{2}\\ &\quad\quad=\frac{1}{2}\left[\frac{1}{4}F_{Q}[\sigma,H_{1}]-I_{\sigma}(H_{1})\right]\\ &\quad\quad\geq 0.\end{split}$ (83)
[…]
Due to Theorem 6 and Theorem 9, if $\varrho$ is pure then
$\tilde{D}_{\rm DPT,decomp}(\varrho,\sigma)^{2}\leq\tilde{D}_{\rm DPT}(\varrho,\sigma)^{2}$ (84)
holds.
For using the quantities based on an optimization over separable states defined above to measure how far a quantum state is from another one, it would be important to show that these quantities fulfill the triangle inequality. Let us now examine this in the case of $\tilde{D}_{\rm DPT,decomp}(\varrho,\sigma).$ We will take advantage of the fact that it is defined as an optimization of an expression given for two pure states and the optimization is taken over all possible separable decompositions of a bipartite state with given marginals.
Theorem 10. The inequality
$\begin{split}&\tilde{D}_{\rm DPT,decomp}(\varrho,\omega)+\tilde{D}_{\rm DPT,decomp}(\omega,\tau)\\ &\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\tilde{D}_{\rm DPT,decomp}(\varrho,\tau)\geq 0,\end{split}$ (85)
where $\varrho,\omega$ and $\tau$ are density matrices, is fulfilled for the case of a pure $\omega.$
Proof. We know that $\tilde{D}_{\rm DPT}(\varrho,\omega)$ fulfills an analogous inequality for $\omega$ being pure or both $\varrho$ and $\tau$ being pure [29], or one of the three states being pure [28]. From this, it follows trivially that Eq. (85) is true for pure states, since for pure states $\tilde{D}_{\rm DPT,decomp}(\varrho,\omega)$=$\tilde{D}_{\rm DPT}(\varrho,\omega)$ holds. However, the triangle inequality can also be proven easily for pure states directly.
Let us assume that $w=|\Omega\rangle\langle\Omega|$ is a pure state. Let us consider the optimal decompositions
$\tilde{D}_{\rm DPT,decomp}(\varrho,\omega)^{2}=\sum_{k}p_{k}\tilde{D}_{\rm DPT}(|\Psi_{k}\rangle,|\Omega\rangle)^{2},$ (86a)
$\tilde{D}_{\rm DPT,decomp}(\omega,\tau)^{2}=\sum_{k}p_{k}\tilde{D}_{\rm DPT}(|\Omega\rangle,|\Upsilon_{k}\rangle)^{2},$ (86b)
$\tilde{D}_{\rm DPT,decomp}(\varrho,\tau)^{2}=\sum_{k}p_{k}\tilde{D}_{\rm DPT}(|\Psi_{k}^{\prime}\rangle,|\Upsilon_{k}^{\prime}\rangle)^{2},$ (86c)
where
$\begin{split}&\varrho=\sum_{k}p_{k}|\Psi_{k}\rangle\langle\Psi_{k}|=\sum_{k}p_{k}|\Psi_{k}^{\prime}\rangle\langle\Psi_{k}^{\prime}|,\\ &\tau=\sum_{k}p_{k}|\Upsilon_{k}\rangle\langle\Upsilon_{k}|=\sum_{k}p_{k}|\Upsilon_{k}^{\prime}\rangle\langle\Upsilon_{k}^{\prime}|.\\ \end{split}$ (87)
In these decompositions states with different indices are not necessarily different, that is why we can use common coefficients $p_{k},$ similarly in the proof of Theorem 1. Then,
$\begin{split}&[\tilde{D}_{\rm DPT,decomp}(\varrho,\omega)+\tilde{D}_{\rm DPT,decomp}(\omega,\tau)]^{2}\\ &\quad\quad=\sum_{k}p_{k}\bigg{[}\tilde{D}_{\rm DPT}(|\Psi_{k}\rangle,|\Omega\rangle)^{2}+\tilde{D}_{\rm DPT}(|\Omega\rangle,|\Upsilon_{k}\rangle)^{2}\bigg{]}\\ &\quad\quad+2\sqrt{\sum_{k}p_{k}\tilde{D}_{\rm DPT}(|\Psi_{k}\rangle,|\Omega\rangle)^{2}}\sqrt{\sum_{k}p_{k}\tilde{D}_{\rm DPT}(|\Omega\rangle,|\Upsilon_{k}\rangle)^{2}}\\ &\quad\quad\geq\sum_{k}p_{k}\bigg{[}\tilde{D}_{\rm DPT}(|\Psi_{k}\rangle,|\Omega\rangle)^{2}+\tilde{D}_{\rm DPT}(|\Omega\rangle,|\Upsilon_{k}\rangle)^{2}\\ &\quad\quad+2\tilde{D}_{\rm DPT}(|\Psi_{k}\rangle,|\Omega\rangle)\tilde{D}_{\rm DPT}(|\Omega\rangle,|\Upsilon_{k}\rangle)\bigg{]}\\ &\quad\quad=\sum_{k}p_{k}[\tilde{D}_{\rm DPT}(|\Psi_{k}\rangle,|\Omega\rangle)+\tilde{D}_{\rm DPT}(|\Omega\rangle,|\Upsilon_{k}\rangle)]^{2}\\ &\quad\quad\geq\sum_{k}p_{k}\tilde{D}_{\rm DPT}(|\Psi_{k}\rangle,|\Upsilon_{k}\rangle)^{2}\\ &\quad\quad\geq\sum_{k}p_{k}\tilde{D}_{\rm DPT}(|\Psi_{k}^{\prime}\rangle,|\Upsilon_{k}^{\prime}\rangle)^{2}=\tilde{D}_{\rm DPT,decomp}(\varrho,\tau)^{2}.\end{split}$ (88)
In the first inequality, we used that for any $x_{k}$ and $y_{k}$ non-negative real numbers and $p_{k}$ probabilities we have the Cauchy-Schwarz inequality
$\begin{split}&\sum_{k}p_{k}x_{k}y_{k}=\sum_{k}(\sqrt{p_{k}}x_{k})(\sqrt{p_{k}}y_{k})\\ &\quad\quad\quad\quad\quad\quad\leq\sqrt{\sum_{k}p_{k}x_{k}^{2}}\sqrt{\sum_{k}p_{k}y_{k}^{2}}.\end{split}$ (89)
[…]
After considering several possibilities of defining the quantum Wasserstein distance based on an optimization over separable states, let us summarize our findings. In many of the examples in Sec. III, we have one of the two possibilities. We had the square of the De Palma-Trevisan distance, $D_{\rm DPT}(\varrho,\sigma)^{2}$ optimized over separable states
$\min_{\{p_{k},|\Psi_{k}\rangle,|\Phi_{k}\rangle\}}\;$ $\sum_{k}p_{k}D_{\rm DPT}(|\Psi_{k}\rangle,|\Phi_{k}\rangle)^{2},$ (90)
where $D_{\rm DPT}(|\Psi_{k}\rangle,|\Phi_{k}\rangle)^{2}=\Delta(|\Psi_{k}\rangle,|\Phi_{k}\rangle)^{2}$ and $\Delta$ is given in Eq. (18). The other possibility was the square of the modified De Palma-Trevisan distance, $\tilde{D}_{\rm DPT}(\varrho,\sigma)^{2}$ optimized over separable states
$\min_{\{p_{k},|\Psi_{k}\rangle,|\Phi_{k}\rangle\}}\;$ $\sum_{k}p_{k}\tilde{D}_{\rm DPT}(|\Psi_{k}\rangle,|\Phi_{k}\rangle)^{2},$ (91)
where $\tilde{D}_{\rm DPT}(|\Psi_{k}\rangle,|\Phi_{k}\rangle)^{2}=E(|\Psi_{k}\rangle,|\Phi_{k}\rangle)^{2}$ and $E$ is given in Eq. (57). We assumed that that the conditions for the marginals given in Eq. (16) are fulfilled in both cases. In the first case, we have a quantity that might have a nonzero self-distance. In the second case, the quantify given by the optimization does not have a self-distance.
IV Quantum $p$-Wasserstein distance based on an optimization over separable states
In this section, we will examine quantum $p$-swap fidelity and quantum $p$-Wasserstein distance definitions based on an optimization over separable states.
Beatty and Stilck França considered the optimization of the trace distance for pure states, and defined quantum $p$-Wasserstein distances with it [40]. In particular, they considered for $p=2$
$\begin{split}&D_{\rm BSF,2}(\varrho,\sigma)^{2}\\ &\quad\quad=\min_{\{p_{k},|\Psi_{k}\rangle,|\Phi_{k}\rangle\}}\;\sum_{k}p_{k}(1-|\langle\Psi_{k}|\Phi_{k}\rangle|^{2}),\end{split}$ (92)
such that the conditions for the marginals given in Eq. (16) are fulfilled. They also defined a quantum $p$-Wasserstein distance as
$\begin{split}&D_{{\rm BSF},p}(\varrho,\sigma)^{2}\\ &\quad\quad=\left[\min_{\{p_{k},|\Psi_{k}\rangle,|\Phi_{k}\rangle\}}\;\sum_{k}p_{k}(1-|\langle\Psi_{k}|\Phi_{k}\rangle|^{2})^{\frac{p}{2}}\right]^{\frac{2}{p}},\end{split}$ (93)
such that for the marginals Eq. (16) are fulfilled. An advantage of the definition is that we can define an expression for a definition given for for pure states and for identical pure states the distance is zero for all $p.$ Moreover, for pure states the distance gives back the underlying metric, which is the trace distance. For a given $p,$ we have just the $p^{th}$ power of the trace distance [40].
Hence, we can see the relation to the quantities we have considered before
$D_{\rm BSF,2}(\varrho,\sigma)^{2}=1-F_{S,\rm sep}(\varrho,\sigma)=2D_{S,\rm sep}(\varrho,\sigma)^{2}.$ (94)
These ideas can be used to calculate $D_{\rm BSF,2}^{2}(\varrho,\sigma)$ numerically based on the methods explained in Refs. [52, 30].
We now arrive at an important theorem.
Theorem 11. If we have a full set of $(d^{2}-1)$ operators $H_{n}$ fulfilling Eq. (70) then
$\begin{split}2D_{\rm BSF,2}(\varrho,\sigma)^{2}&=\tilde{D}_{\rm DPT,sep}(\varrho,\sigma)^{2}\\ &=D_{\rm DPT,sep}(\varrho,\sigma)^{2}-2(d-1).\end{split}$ (95)
holds.
[…]
Thus, the definitions of $D_{\rm BSF,2}(\varrho,\sigma)^{2}$ given in Ref. [40] and $D_{\rm DPT,sep}(\varrho,\sigma)^{2}$ given in Ref. [30] are unexpectedly linear functions of each other.
We can summarize some relevant relations bounding $D_{\rm BFS,2}(\varrho,\sigma)^{2}$ from Eqs. (6), (37), (43), and (94) as
$\begin{split}&D_{\rm Bures}(\varrho,\sigma)^{2}/2=1-\sqrt{F(\varrho,\sigma)}\\ &\quad\quad\quad\quad\leq 2D_{S}(\varrho,\sigma)^{2}\\ &\quad\quad\quad\quad\leq D_{\rm BFS,2}(\varrho,\sigma)^{2}=2D_{S,\rm sep}(\varrho,\sigma)^{2}\\ &\quad\quad\quad\quad\leq 1-F(\varrho,\sigma).\end{split}$ (96)
Note that for qubits the last inequality is saturated due to Theorem 4. Thus, for qubits
$D_{\rm BFS,2}(\varrho,\sigma)^{2}=1-F(\varrho,\sigma)$ (97)
holds. The last inequality in Eq. (96) is also saturated and Eq. (97) holds, if one of the two states is pure due to Eq. (80).
Next we will define an optimization of the powers of the overlap knowing that an optimization of some power of an expectation value over a quantum state can be expressed an an optimization over a separable state with a given marginal [52]. Let us define the $p$-swap fidelity based on an optimization over separable states as
$F^{(p)}_{S,{\rm sep}}(\varrho,\sigma)=\max_{\{p_{k},\Phi_{k},\Psi_{k}\}}\sum_{k}p_{k}|\langle\Phi_{k}|\Psi_{k}\rangle|^{2p},$ (98)
which can also be rewritten as
$F^{(p)}_{S,{\rm sep}}(\varrho,\sigma)=\max_{\varrho_{123...2p}\in\mathrm{Sep}_{p}}{\rm Tr}(\varrho_{123..2p}S_{p}),$ (99)
[…]
where $P_{\rm sym}$ is the projector to the symmetric subspace of the quantum states of $p$ particles, and $𝟙\openone$ acts also on $p$ particles. It is instructive to write down the $\varrho_{123...2p}$ consistent with these conditions as
$\sum_{k}p_{k}(|\Psi_{k}\rangle\langle\Psi_{k}|)^{\otimes p}\otimes(|\Phi_{k}\rangle\langle\Phi_{k}|)^{\otimes p}.$ (101)
The value of the maximum will remain the same, even if the conditions with the projector to the symmetric subspace are neglected (see Observations 1 and 4 in Ref. [30] for a similar problem), see also Appendix B.
The quantum $p$-Wasserstein distance can be calculated based on similar ideas. We find that
$D_{{\rm BSF},p}(\varrho,\sigma)^{2}=\left[\min_{\varrho_{123...2p}\in\mathrm{Sep}_{p}}{\rm Tr}\left(\varrho_{123..2p}O\right)\right]^{\frac{2}{p}},$ (102)
where the operator for the optimization is
[…]
and the conditions for the optimization are the same. Here, $S^{(k,l)}$ is the swap operator acting between particle $k$ and $l.$ For pure states, we have
$\begin{split}&{\rm Tr}(|\Phi\rangle\langle\Phi|^{\otimes p}|\Psi\rangle\langle\Psi|^{\otimes p}O)\\ &\quad\quad\quad\quad\quad\propto\left[\sum_{n=1}^{d^{2}-1}(\langle{H_{n}}\rangle_{\Psi}-\langle{H_{n}}\rangle_{\Phi})^{2}\right]^{\frac{p}{2}},\end{split}$ (104)
where $O$ acts on a $2^{p}\times 2^{p}$ bipartite state, and $H_{n}$ is a full set of $(d^{2}-1)$ operators fulfilling Eq. (70) mentioned before. These quantities can be calculated numerically based on the methods explained in Refs. [52, 30], e. g., by approximating the separable states using the separability criterion based on symmetric extensions [61, 62, 63]. We add that quantum $p$-Wasserstein distances have been considered for the case of optimization over all quantum states, rather than separable states [28].
Let us summarize our findings for the case of if we consider a full set of $H_{n}$ matrices. Then, several of the quantities considered in this paper are equal to each other
$\begin{split}&D_{\rm DPT,sep}(\varrho,\sigma)^{2}-2(d-1)=\tilde{D}_{\rm DPT,sep}(\varrho,\sigma)^{2}\\ &\quad\quad=\tilde{D}_{\rm DPT,decomp}(\varrho,\sigma)^{2}=4D_{\rm S,sep}(\varrho,\sigma)^{2}\\ &\quad\quad=2(1-F_{\rm S,sep}(\varrho,\sigma)^{2})=2D_{\rm BSF,2}(\varrho,\sigma)^{2},\end{split}$ (105)
which is the consequence of Theorem 8, and Eqs. (74), (76), and (94). Thus, in Theorem 10 we proved the triangle inequality also for all the quantities above.
Conclusions
We studied several possible definitions of the quantum Wasserstein distance. We considered some that are based on an optimization over general bipartite states. Then, we considered reducing the optimization to separable states, and determined the corresponding quantities. We studied the relations between these quantities. We proved a triangle inequality for some of these quantities for the case when of the states is pure. The details of the numerical methods we have been using are in Appendix C.
Acknowledgements.
We thank I. Apellaniz, M. Eckstein, F. Fröwis, I. L. Egusquiza, C. Klempt, J. Kołodyński, M. W. Mitchell, M. Mosonyi, G. Muga, J. Siewert, Sz. Szalay, K. Życzkowski, T. Vértesi, G. Vitagliano, and D. Virosztek for discussions. We acknowledge the support of the EU (QuantERA MENTA, QuantERA QuSiED, COST Action CA23115), the Spanish MCIU (Grant No. PCI2022-132947), the Basque Government (Grant No. IT1470-22), and the National Research, Development and Innovation Office of Hungary (NKFIH) (2019-2.1.7-ERA-NET-2021-00036). We thank the National Research, Development and Innovation Office of Hungary (NKFIH) within the Quantum Information National Laboratory of Hungary. We acknowledge the support of the Grant No. PID2021-126273NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. We thank the “Frontline” Research Excellence Programme of the NKFIH (Grant No. KKP133827). We thank Project no. TKP2021-NVA-04, which has been implemented with the support provided by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021-NVA funding scheme. G. T. is thankful for a Bessel Research Award from the Humboldt Foundation.
Appendix A Proof of Theorem 8
We show that the inequality in Eq. (67) is saturated if we consider a full set of $(d^{2}-1)$ operators $H_{n}$ fulfilling Eq. (70).
In this case, we need to use that for all pure states we the equality in Eq. (75) holds. Hence for all $\{p_{k},\Psi_{k}\}$ decompositions we have
$\sum_{k}p_{k}\sum_{n=1}^{d^{2}-1}(\Delta H_{n})^{2}_{\Psi_{k}}=2(d-1).$ (106)
Similarly, for all $\{p_{k},\Phi_{k}\}$ decompositions we have
$\sum_{k}p_{k}\sum_{n=1}^{d^{2}-1}(\Delta H_{n})^{2}_{\Phi_{k}}=2(d-1).$ (107)
Hence, for the self-distance
$D_{\rm DPT,sep}(\varrho,\varrho)^{2}=D_{\rm DPT,sep}(\sigma,\sigma)^{2}=2(d-1)$ (108)
holds. Thus, the self-distances are constants independent from the quantum state. The decomposition $\{p_{k},\Psi_{k},\Phi_{k}\}$ that minimizes
$\min_{\{p_{k},|\Psi_{k}\rangle,|\Phi_{k}\rangle\}}\;(\langle{H_{n}}\rangle_{\Psi_{k}}-\langle{H_{n}}\rangle_{\Phi_{k}})^{2},$ (109)
will saturate the inequality in Eq. (67).
We show a simple example with $N=1$ where the two sides in Eq. (67) are not equal to each other. For $N=1$, Eq. (67) can be rewritten as Eq. (68). Let us now consider $H_{1}=\sigma_{z}$ and
[…]
For these, for the quantum Fisher information
$F_{Q}[\varrho,H_{1}]=F_{Q}[\sigma,H_{1}]=0$ (111)
hold, and
$D_{\rm DPT,sep}(\varrho,\sigma)^{2}=\tilde{D}_{\rm DPT,sep}(\varrho,\sigma)^{2}$ $=1,$ (112a)
$\tilde{D}_{\rm DPT,decomp}(\varrho,\sigma)^{2}$ $=1/2.$ (112b)
An optimal coupling for Eq. (112a) is
$\varrho\otimes\sigma=\frac{1}{2}(|0\rangle\langle 0|\otimes|0\rangle\langle 0|+|1\rangle\langle 1|\otimes|0\rangle\langle 0|).$ (113)
For calculating Eq. (112b), we need the decomposition
$\varrho\otimes\sigma=\frac{1}{2}(|0\rangle\langle 0|_{x}\otimes|0\rangle\langle 0|+|1\rangle\langle 1|_{x}\otimes|0\rangle\langle 0|),$ (114)
where |.$|.\rangle_{x}$ refers to the $\sigma_{x}$ basis.
Appendix B Symmetry properties of the optimal state
In this Appendix, we prove that there is an optimal state for the maximization in Eq. (99) that is of the form Eq. (101).
From the inequality between the arithmetic and geometric mean follows that
$\frac{1}{p}\sum_{n=1}^{p}{\rm Tr}(A_{n})\geq{\rm Tr}\left(\bigotimes_{n=1}^{p}A_{n}\right)^{\frac{1}{p}},$ (115)
where $A_{n}$ are positive-semidefinite Hermitian matrices, which can be rewritten as
${\rm Tr}\left[\left(\frac{1}{p}\sum_{n=1}^{p}A_{n}\right)^{\otimes p}\right]\geq{\rm Tr}\left(A_{1}\otimes A_{2}\otimes...\otimes A_{p}\right).$ (116)
We will also use that for $A$ and $B$ matrices Eq. (40) holds. Then, we can write an inequality relevant for our problem as
$\begin{split}&{\rm Tr}\left(\sum_{k}{\tilde{p}_{k}}\bigotimes_{n=1}^{p}\varrho_{k}^{(n,n+p)}S_{p}\right)\\ &\quad\quad\quad\quad\geq{\rm Tr}\left(\sum_{k}{\tilde{p}_{k}}\bigotimes_{n=1}^{p}\varrho^{(n,n+p)}_{k,n}S_{p}\right),\end{split}$ (117)
where we define the two-particle product density matrices as
$\varrho_{k,n}=|\Psi_{k}^{(n)}\rangle\langle\Psi_{k}^{(n)}|\otimes|\Phi_{k}^{(n)}\rangle\langle\Phi_{k}^{(n)}|,$ (118)
where for all $n=1,2,...,p$ we have
$\varrho$ $=\sum_{k}{\tilde{p}_{k}}|\Psi_{k}^{(n)}\rangle\langle\Psi_{k}^{(n)}|,$ (119a)
$\sigma$ $=\sum_{k}{\tilde{p}_{k}}|\Phi_{k}^{(n)}\rangle\langle\Phi_{k}^{(n)}|,$ (119b)
and the average two-particle separable density matrix is
$\varrho_{k}=\frac{1}{p}\sum_{n=1}^{p}\varrho_{k,n}.$ (120)
The superscript $(n,n+p)$ indicates that the two-particle density matrix corresponds to spins $n$ and $(n+p).$ Let us consider the state
$\varrho_{12}=\sum_{k}{\tilde{p}_{k}}\varrho_{k},$ (121)
which can be written as in Eq. (15). Note that the reduced states are the following
${\rm Tr}_{2}\left(\varrho_{12}\right)=\varrho,\quad{\rm Tr}_{1}\left(\varrho_{12}\right)=\sigma.$ (122)
The inequality in Eq. (117) is saturated if
$\varrho_{k,n}=\varrho_{k,1}$ (123)
for all $k$ and $n.$
Now we have to use that
${\rm Tr}\left(\bigotimes_{n=1}^{p}\varrho_{k}^{(n,n+p)}S_{p}\right)={\rm Tr}(\varrho_{k}S)^{p}$ (124)
is a convex function of the two-body density matrix $\varrho.$ This statement is true for general quantum states for an even $p,$ however, it is also true for an odd $p$ for separable states due to the fact that for a separable state Eq. (42) holds. Thus, when maximized over a convex set, the expression in Eq. (124) takes its maximum at the extreme points of the set.
Hence, we can see that that there is a $\{p_{k},\tau_{k}\}$ decomposition such that
$\begin{split}&{\rm Tr}\left(\sum_{k}{p_{k}}\bigotimes_{n=1}^{p}\tau_{k}^{(n,n+p)}S_{p}\right)\\ &\quad\quad\quad\quad\geq{\rm Tr}\left(\sum_{k}\tilde{p}_{k}\bigotimes_{n=1}^{p}\varrho_{k}^{(n,n+p)}S_{p}\right),\end{split}$ (125)
where the two-particle pure state is given as
$\tau_{k}=|\Psi_{k}\rangle\langle\Psi_{k}|\otimes|\Phi_{k}\rangle\langle\Phi_{k}|,$ (126)
and $\varrho$ and $\sigma$ is decomposed as in Eq. (16). Together with Eq. (117) this indicates that there is an optimal state of the form Eq. (101). The quantum state appearing on the left-hand side of Eq. (125) is of that form.
Appendix C Details of the numerical calculations
We used MATLAB [64] for numerical calculations. We used the semidefinite solver MOSEK [65] and the front-end YALMIP [66]. We also used the QUBIT4MATLAB package [67, 68]. $D_{\rm DPT}(\varrho,\sigma)^{2}$ given in Eq. (1) can be obtained using semidefinite programming. $D_{\rm DPT,sep}(\varrho,\sigma)^{2}$ in Eq. (12) need an optimization over separable states. Instead of an optmization over separable states, we can carry out numerically an optimization over the set of PPT states (i. e., states with a positive partial transpose). Since it is a superset of separable states, if we maximize, we obtain an upper bound on the maximum over separable states, and if we minimize, we obtain a lower bound on the minimum over separable states. However, for two qubits we will get the true optimal value, since in that case the set of PPT states equals the set of separable states. For a detailed discussion about how to compute numerically the quantum Wasserstein distance via an optimization over separable states see Ref. [30].
We have already included the routines computing the various quantum Wasserstein distance measures for Ref. [30], such as wdistsquare_GMPC.m, wdistsquare_DPT_ppt.m, and wdistsquare_DPT.m. We included the various types of the variance-like quantities as wvar_GMPC_ppt.m, wvar_GMPC.m, wvar_DPT_ppt.m, and wvar_DPT.m. The usage of these routines is demonstrated in example_wdistsquare.m.
For this publication, we added swapfidelity.m, swapfidelity_ppt.m, and example_swapfidelity.m. Instead of the command swapfidelity_ppt.m, the old command wdistsquare_DPT_ppt.m could also be used with an appropriate choice of the $H_{n}$ operators. This is due to the relation given in Eq. (105) for the case when we consider a full set of $(d^{2}-1)$ traceless operators $H_{n},$ which are the SU(d) generators. We included the new command for convenience.